OpenStudy (anonymous):

Differentiate 2/((1-x^2)^2) What I'm having trouble with is when you move (1-x^2)^2 to the numerator is it all negative? I need help on the first step only, i think i can do the rest.

ganeshie8 (ganeshie8):

use the exponents property : \(\large \frac{1}{a^{n}} = a^{-n}\)

OpenStudy (anonymous):

\( \dfrac{2}{(1-x^2)^2}\)

ganeshie8 (ganeshie8):

\(\large \frac{2}{(1-x^2)^2} = 2(1-x^2)^{-2}\)

OpenStudy (anonymous):

Ohhhh so only the power again...

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

I keep getting that mixed up for some reason.

OpenStudy (anonymous):

quotient rule is just alot of movement going on and can easily be viewed as the simpler to work product rule

OpenStudy (anonymous):

We haven't learnt the quotient rule yet, and for this question the chain rule is needed.

OpenStudy (anonymous):

the chain rule is always needed regardless of the rule :) it just that in most cases an x' = dx/dx = 1 pops out